OpenStudy (anonymous):

What is the equation of a line parallel to y= -1/4x +7 going through the point ( 8,27) ? Help please

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (skullpatrol):

Hint: what do parallel lines have in common?

OpenStudy (anonymous):

They have the same slope right?

OpenStudy (skullpatrol):

Correct :-) In the equation form y = mx + b Which letter stands for the slope?

OpenStudy (anonymous):

The m

OpenStudy (skullpatrol):

Correct again! So we know the answer will have the form y = (-1/4)x + b And we are given the point (x, y) on the line as (8, 27) Now just substitute those values in the equation form and solve for b. Like this y = (-1/4)x + b 27 =(-1/4)8 + b b = 29 This gives the answer as y = (-1/4)x + 29.

OpenStudy (skullpatrol):

Any questions?

OpenStudy (jack1):

nice work skulls